OpenStudy (anonymous):

Generate the first 5 terms of this sequence: f(1) = 2 and f(2) = 3, f(n) = f(n - 1) + f(n - 2), for n > 2. 2, 3, 5, 7, 9 2, 3, 4, 7, 11 2, 3, 5, 9, 11 2, 3, 5, 8, 13

OpenStudy (anonymous):

f(3) = f(2) + f(1) => 3 + 2 = 5 f(4) = f(3) + f(2) => 5 + 3 = 8 f(5) = f(4) + f(3) => 8 + 5 = 13 Therefore the sequence is 2, 3, 5, 8, 13,....

OpenStudy (anonymous):

thanks!